# Isoperimetric inequalities and sharp upper bounds for Aharonov-Bohm eigenvalues on surfaces

### 13 Apr 2026 v1

@@ 3.2. Monotonicity of the function $G$
where $\beta$ is the inverse of $\alpha$. Now, if $\Omega$ is a simply connected Riemannian surface with boundary of area $\mathcal{A}$, boundary length $\ell$, and gaussian curvature bounded by $1$, we have the following isoperimetric inequality:
@@ 3.3. Optimization of a 1-d problem
We have now to consider the following number
@@
where $\mathcal{F}=\{f\in L^{2}(0,M):\sqrt{G}f^{\prime},f/\sqrt{G}\in L^{2}(0,M)\}$, where now $G:(0,M)\to(0,\infty)$ is any smooth, positive function satisfying
@@
Note that $\kappa_{1}(\Omega,A^{(\nu)}_{z_{0}})=\kappa_{1}(G_{z_{0}})$ where $G_{z_{0}}$ has been defined in (3.5). Recall also that $G^{\star}=a(4\pi-a)$. We want to optimize $\kappa_{1}(G)$ with respect the weight $G$. In particular we have the following lemma
@@
We consider then $\kappa_{1}(G_{t})$ and compute its derivative with respect to $t$. Since the eigenvalue is simple, we can apply Feynman-Hellmann formula [11, VII-§4, p.408, n. 4.56] to get
@@
where $f_{t}$ is the first $L^{2}$-normalized eigenfunction corresponding to the eigenvalue $\kappa_{1}(G_{t})$, chosen to be positive. In order to conclude we have to prove that
@@
Second step. It remains to consider the case in which there exists $a_{0}\in(0,M)$ such that $R(a_{0})\leq-2\pi\nu$. This implies, by (3.12), that $R^{\prime}<0$ on $(a_{0},M)$, and hence $R(M)<-2\pi\nu$, which is a contradiction with the fact that $R(M)=0$. ∎
@@ 3.4. Conclusion of the proof
The proof of Theorem 2.1 follows now from (3.2), (3.3) and Lemmas 3.2 and 3.3:
@@
4. Proof of Theorems 2.3 and 2.5
The proof is similar to that of Theorem 2.1. However, note that Theorems 2.3 and 2.5 give an upper bound which is not isoperimetric. In particular, in the case of a closed surface $\Sigma$, the bounded gaussian curvature hypothesis implies $|\Sigma|\geq|\mathbb{S}^{2}_{K}|$, and if $\Sigma\neq\mathbb{S}^{2}_{K}$, the inequality is strict. As in Section 3, we may suppose that $K=1$. We will prove Theorem 2.3 first.
4.1. Reduction to 1-d problem
@@
(4.4) $w_{1}(\Sigma,A)=\min_{0\neq f\in\mathcal{F}}\frac{\int_{0}^{M}\left(G(a)f^{\prime}(a)^{2}+\frac{4pi^{2}\nu^{2}}{G(a)}f(a)^{2}\right)da}{\int_{0}^{M}f(a)^{2}da},$
@@ 4.2. Optimization of $1$-d problem
The proof is identical to that of Lemma 3.3. The only difference from the proof of Lemma 3.3 lies in the second step, namely the treatment of the case $R\leq-2\pi\nu$ since now also $a=4\pi$ is a singular point.
Second step. Assume by contradiction that there exists a point $a_{0}\in(0,M)$ such that $R(a_{0})\leq-2\pi\nu$. We see from (3.12) that $R^{\prime}(a_{0})<0$, and this implies, still using (3.12), that $R$ is decreasing in the whole interval $(a_{0},4\pi)$. Hence limit $\lim_{a\rightarrow 4\pi^{-}}R(a)$ exists, so we have two cases
Case $1$: $\lim_{a\rightarrow 4\pi^{-}}R(a)$ is finite. In particular then, $\lim_{a\rightarrow 4\pi^{-}}R(a)<-2\pi\nu$. We integrate (3.12) and we obtain
@@
But now we recall that $G_{t}(a)\sim 4\pi(4\pi-a)$ when $a\rightarrow 4\pi^{-}$, hence, since $\lim_{a\rightarrow 4\pi^{-}}R(a)<-2\pi\nu$, we get that the right-hand side of the above equation is unbounded as $a\to 4\pi^{-}$, while the left-hand side remains bounded.
Case $2$: $\lim_{a\rightarrow 4\pi^{-}}R(a)=$-$\infty$. From (3.12), recalling that $G_{t}(a)\sim 4\pi(4\pi-a)$ as $a\to 4\pi^{-}$, we obtain
$\lim_{a\rightarrow 4\pi^{-}}\frac{(4\pi-a)R^{\prime}}{R^{2}}=-\frac{1}{4\pi}.$
@@
$\frac{1}{(4\pi-a)}\left(-\frac{1}{4\pi}-\epsilon\right)\leq\left(-\frac{1}{R}\right)^{\prime}\leq\frac{1}{(4\pi-a)}\left(-\frac{1}{4\pi}+\epsilon\right)$
@@
$\left(-\frac{1}{4\pi}-\epsilon\right)\ln\left({\frac{4\pi-\bar{a}}{4\pi-a}}\right)\leq\frac{1}{R(a)}-\frac{1}{R(\bar{a})}\leq\left(-\frac{1}{4\pi}+\epsilon\right)\ln\left({\frac{4\pi-\bar{a}}{4\pi-a}}\right)$
In particular from the above inequality we can conclude that $\lim_{a\rightarrow 4\pi^{-}}R(a)=0$ that is a contradiction.
The conclusion now is the same as in Lemma 3.3. ∎
4.3. A sequence of weight functions
Due to the volume bound $|\Sigma|\geq|\mathbb{S}^{2}|$, we cannot directly compare $w_{1}(G)$ with $w_{1}(G^{\star})$, simply because $G,G^{\star}$ are defined in different intervals: $(0,M)$ and $(0,4\pi)$, with $4\pi\leq M$ (recall, $M=|\Sigma|$). Moreover $M=4\pi$ if and only if $\Sigma=|\mathbb{S}^{2}|$.
We define a sequence of weights $G_{\varepsilon}$ as follows. Let $\varepsilon_{0}>0$ be a sufficiently small number to be fixed later, and let $\varepsilon\in(0,\varepsilon_{0})$. Let
@@ 4.3. A sequence of weight functions
and $G>0$ in $(0,M)$. Then there exists $\varepsilon_{0}>0$ such that for all $\varepsilon\in(0,\varepsilon_{0})$, we have $G\geq G_{\varepsilon}$, where
$G_{\varepsilon}=\begin{cases}G^{\star}\,,&{\rm in\ }(0,a_{\varepsilon})\\ \varepsilon\in(a_{\varepsilon},M-b_{\varepsilon})\,,&{\rm in\ }(a_{\varepsilon},M-b_{\varepsilon})\\ G^{\star}(a-M)\,,&{\rm in\ }(M-b_{\varepsilon},M).\end{cases}$
See Figure 1. The function $G_{\varepsilon}$ is continuous and uniformly converges to $G^{\star}$ in $(0,4\pi)$ and to 0 in $(4\pi,M)$.
Figure 1. the function $G_{\varepsilon}$
@@
From Lemma 4.3 we deduce that, for all $\varepsilon\in(0,\varepsilon_{0})$
(4.7) $w_{1}(G)\leq w_{1}(G_{\varepsilon}).$
Now, the upper bound (2.1) in Theorem 2.3 is a consequence of the following lemma and from the fact that $\mu_{1}(\mathbb{S}^{2},A^{\star})=\min_{n\in\mathbb{Z}}|n-\nu|$, which equals $\nu(\nu+1)$ if $\nu\in(0,\frac{1}{2}]$. The explicit expression of $\mu_{1}(\mathbb{S}^{2},A^{\star})$ is proved in Appendix A. Recall that by $\lambda_{1}(\Omega,A)$ we denote the first magnetic Dirichlet eigenvalue on $\Omega$ with potential $A$, see (1.1).
@@
    $w_{1}(G_{\varepsilon})\leq\lambda_{1}(D_{a_{\varepsilon}},A^{\star})$, where $\lambda_{1}(D_{a_{\varepsilon}},A^{\star})$ is the first magnetic Dirichlet eigenvalue of a geodesic disk $D_{a_{\varepsilon}}\subset\mathbb{S}^{2}$ of area $a_{\varepsilon}$ with center $x_{0}$ and with $A^{\star}=-2\pi\nu\star d\psi_{x_{0}}=\nu d\theta$
@@
(4.8) $w_{1}(G_{\varepsilon})=\min_{0\neq f\in\mathcal{F}}\frac{\int_{0}^{M}\left(G_{\varepsilon}(a)f^{\prime}(a)^{2}+\frac{4\pi^{2}\nu^{2}}{G_{\varepsilon}(a)}f(a)^{2}\right)da}{\int_{0}^{M}f^{2}(a)da}\\ \leq\min_{\begin{subarray}{c}0\neq f\in\mathcal{F}\\ f\equiv 0{\rm\ on\ }(a_{\varepsilon},M)\end{subarray}}\frac{\int_{0}^{M}\left(G_{\varepsilon}(a)f^{\prime}(a)^{2}+\frac{4\pi^{2}\nu^{2}}{G_{\varepsilon}(a)}f(a)^{2}\right)da}{\int_{0}^{M}f^{2}(a)da}\\ =\min_{\begin{subarray}{c}0\neq f\in\mathcal{F}\\ f(a_{\varepsilon})=0\end{subarray}}\frac{\int_{0}^{a_{\varepsilon}}\left(G^{\star}(a)f^{\prime}(a)^{2}+\frac{4\pi^{2}\nu^{2}}{G^{\star}(a)}f(a)^{2}\right)da}{\int_{0}^{a_{\varepsilon}}f^{2}(a)da}=\lambda_{1}(D_{a_{\varepsilon}},A^{\star}).$
The last equality holds since for the disk with Aharonov-Bohm potential with pole in the center and flux $\nu\in(0,\frac{1}{2}]$, there exists a first real and radial eigenfunction, as in the Neumann case. This proves i). As for ii), we will prove in Lemma A.4 that the first magnetic Dirichlet eigenvalue on a spherical cap with pole at its center coverges to the first eigenvalue on the sphere with two opposite poles when the cap invades the whole sphere. ∎
@@
(4.9) $\mu_{1}(\mathbb{S}^{2},A^{\star})=\mu_{1}(\Sigma,A)\leq w_{1}(G)\leq w_{1}(G_{\varepsilon_{1}})<w_{1}(G_{\varepsilon_{2}})\\ <\cdots<w_{1}(G_{\varepsilon_{n}})\leq\lambda_{1}(D_{a_{\varepsilon_{n}}},A^{\star})\to\mu_{1}(\mathbb{S}^{2},A^{\star}).$
By Lemma 4.3 we have that the inequalities $w_{1}(G_{\varepsilon_{n}})<w_{1}(G_{\varepsilon_{n+1}})$ are strict, since by contruction $G_{\varepsilon_{n}}>G_{\varepsilon_{n+1}}$ on a set of positive measure. Hence all the inequalities must be equalities, which is not possible, unless $M=4\pi$ and hence $\Sigma=\mathbb{S}^{2}$. Thus we are in the case of the sphere $\mathbb{S}^{2}$ with two poles $x_{0},y$ as in Corollary 2.4. If equality holds, then $G=G^{\star}$ in $(0,4\pi)$ and this implies that $y=-x_{0}$ by Lemma 3.2.
@@
The proof of Theorem 2.5 is exactly the same as that of Theorem 2.3, except that now $G_{\varepsilon}$ is defined as
$G_{\varepsilon}=\begin{cases}G^{\star}\,,&{\rm in\ }(0,a_{\varepsilon})\\ \varepsilon\in(a_{\varepsilon},M)\,,&{\rm in\ }(a_{\varepsilon},M).\end{cases}$
for all positive $\varepsilon\leq\varepsilon_{0}$, for a suitable $\varepsilon_{0}>0$. The equality case is discussed as for Theorem 2.3.
@@
Let $\mathcal{C}=\{(x,y,z)\in\mathbb{R}^{3}:x^{2}+y^{2}=1,0\leq z\leq L\}$ where $L>0$. Then we consider $\Omega_{+}:=\{(x,y,z)\in\mathbb{R}^{3}:x^{2}+y^{2}+(z-L)^{2}=1\,,z>L\}$ and $\Omega_{-}:=\{(x,y,z)\in\mathbb{R}^{3}:x^{2}+y^{2}+z^{2}=1\,,z<0\}$. Set $\Omega=\overline{\mathcal{C}\cup\Omega_{1}\cup\Omega_{2}}$ and consider on the hypersurface $\Omega$ the induced metric from the ambient Euclidean space. Note that the metric on $\Omega$ is just of class $C^{1,1}$, but it is enough for our purposes (one can smooth out the metric near the junctures). We also note that the Gaussian curvature belongs $L^{\infty}(\Omega)$, and equals 0 in $\mathcal{C}$, while it equals $1$ in $\Omega_{\pm}$. Let $\mu_{1}(\Omega,A)$ denote the first Aharonov-Bohm eigenvalue on $\Omega$ with magnetic potential $A$, where $A$ is the restriction to $\Omega$ of the $1$-form $\nu d\theta$, where $\theta$ denotes the polar coordinate in the $x_{1}x_{2}$-plane. $A$ has flux $\nu$ around any loop containing the points of coordinates $p_{+}=(0,0,L+1)$ and $p_{-}=(0,0,-1)$.
In Appendix A we prove that $\mu_{1}(\Omega_{\pm},A)=\nu(\nu+1)$ when $\nu\in(0,\frac{1}{2}]$, while in Appendix B we prove that $\mu_{1}(\mathcal{C},A)=\nu^{2}$.
Let $u$ to be an eigenfunction on $\Omega$ associated with $\mu_{1}(\Omega,A)$. From the variational characterization of the eigenvalues we have
$\nu^{2}\int_{\mathcal{C}}|u|^{2}$ $\leq\int_{\mathcal{C}}|d^{A}u|^{2}$
$\nu^{2}\int_{\Omega_{\pm}}|u|^{2}\leq$ $\nu(\nu+1)\int_{\Omega_{\pm}}|u|^{2}\leq\int_{\Omega_{\pm}}|d^{A}u|^{2}$
Summing the above inequalities we obtain
$\nu^{2}\int_{\Omega}|u|^{2}\leq\int_{\Omega}|d^{A}u|^{2},$
and from the variational characterization of $\mu_{1}(\Omega,A)$ we get $\nu^{2}\leq\mu_{1}(\Omega,A)$.
Now, rescaling the metric by a factor $\frac{1}{\sqrt{K}}$, $K>1$, we have that the gaussian curvature of $\Omega_{K}\doteq\frac{1}{\sqrt{K}}\Omega$ is bounded above by $K$ and that $\mu_{1}(\Omega,A)\geq\nu^{2}K$. Note that this bound does not depend on the choice of $L$, and since $|\Omega_{K}|=\frac{2\pi(L+2)}{K}$ we can always choose $L$ such that $|\Omega_{k}|=|\mathbb{S}^{2}|=4\pi$.
4.8. Sharpness of Theorem 2.5
It is enough to consider the example in the previous subsection and remove one of the two spherical caps $\Omega_{+}$ or $\Omega_{-}$.
Appendix A Eigenvalues on the sphere

### 29 Apr 2026 v2

@@ 3.2. Monotonicity of the function $G$
where $\beta$ is the inverse of $\alpha$. Now, if $\Omega$ is a simply connected Riemannian surface with boundary of area $\mathcal{A}$, boundary length $\ell$, and gaussian curvature bounded above by $1$, we have the following isoperimetric inequality:
@@ 3.3. Optimization of a 1-d problem
Next we consider the following number
@@
where $\mathcal{F}=\{f\in L^{2}(0,M):\sqrt{G}f^{\prime},f/\sqrt{G}\in L^{2}(0,M)\}$ and $G:(0,M)\to(0,\infty)$ is any smooth, positive function satisfying
@@
Note that $\kappa_{1}(\Omega,A^{(\nu)}_{z_{0}})=\kappa_{1}(G_{z_{0}})$ where $G_{z_{0}}$ has been defined in (3.5). Recall also that $G^{\star}=a(4\pi-a)$. We want to optimize $\kappa_{1}(G)$ with respect to the weight $G$. In particular we have the following lemma.
@@
We compute the derivative of $\kappa_{1}(G_{t})$ with respect to $t$. Since the eigenvalue is simple, we can apply Feynman-Hellmann formula [11, VII-§4, p.408, n. 4.56] to get
@@
where $f_{t}$ is the first positive $L^{2}$-normalized eigenfunction corresponding to the eigenvalue $\kappa_{1}(G_{t})$. In order to conclude we have to prove that
@@
Second step. It remains to consider the case in which there exists $a_{0}\in(0,M)$ such that $R(a_{0})\leq-2\pi\nu$. This implies, by (3.12), that $R^{\prime}<0$ on $(a_{0},M)$, and hence $R(M)<-2\pi\nu$, which is a contradiction with $R(M)=0$. ∎
@@ 3.4. Conclusion of the proof
The proof of Theorem 2.1 now follows from (3.2), (3.3) and Lemmas 3.2 and 3.3:
@@
4. Proof of Theorems 2.4 and 2.3
The proof is similar to that of Theorem 2.1. However, note that Theorems 2.4 and 2.3 give an upper bound which is not isoperimetric. In particular, in the case of a closed surface $\Sigma$, the bound on the gaussian curvature implies that $|\Sigma|\geq|\mathbb{S}^{2}_{K}|$, and if $\Sigma\neq\mathbb{S}^{2}_{K}$, the inequality is strict. As in Section 3, we may suppose that $K=1$. We will prove Theorem 2.4 first.
4.1. Proof of Theorem 2.4: reduction to a 1-d problem
@@
(4.4) $w_{1}(\Sigma,A)=\min_{0\neq f\in\mathcal{F}}\frac{\int_{0}^{M}\left(G(a)f^{\prime}(a)^{2}+\frac{4\pi^{2}\nu^{2}}{G(a)}f(a)^{2}\right)da}{\int_{0}^{M}f(a)^{2}da},$
@@ 4.2. Optimization of the $1$-d problem
The proof is identical to that of Lemma 3.3. The only difference from the proof of Lemma 3.3 lies in the second step, namely the treatment of the case $R\leq-2\pi\nu$ since now also $a=M$ is a singular point.
Second step. Assume by contradiction that there exists a point $a_{0}\in(0,M)$ such that $R(a_{0})\leq-2\pi\nu$. We see from (3.12) that $R^{\prime}(a_{0})<0$, and this implies, still using (3.12), that $R$ is decreasing in the whole interval $(a_{0},M)$. Hence limit $\lim_{a\rightarrow M^{-}}R(a)$ exists, so we have two cases
Case $1$: $\lim_{a\rightarrow M^{-}}R(a)$ is finite. In particular then, $\lim_{a\rightarrow M^{-}}R(a)<-2\pi\nu$. We integrate (3.12) and we obtain
@@
But now we recall that $G_{t}(a)\sim 4\pi(M-a)$ when $a\rightarrow M^{-}$, hence, since $\lim_{a\rightarrow M^{-}}R(a)<-2\pi\nu$, we get that the right-hand side of the above equation is unbounded as $a\to M^{-}$, while the left-hand side remains bounded.
Case $2$: $\lim_{a\rightarrow M^{-}}R(a)=$-$\infty$. From (3.12), recalling that $G_{t}(a)\sim 4\pi(M-a)$ as $a\to M^{-}$, we obtain
$\lim_{a\rightarrow M^{-}}\frac{(M-a)R^{\prime}}{R^{2}}=-\frac{1}{4\pi}.$
@@
$\frac{1}{(M-a)}\left(-\frac{1}{4\pi}-\epsilon\right)\leq\left(-\frac{1}{R}\right)^{\prime}\leq\frac{1}{(M-a)}\left(-\frac{1}{4\pi}+\epsilon\right)$
@@
$\left(-\frac{1}{4\pi}-\epsilon\right)\ln\left({\frac{M-\bar{a}}{M-a}}\right)\leq\frac{1}{R(a)}-\frac{1}{R(\bar{a})}\leq\left(-\frac{1}{4\pi}+\epsilon\right)\ln\left({\frac{M-\bar{a}}{M-a}}\right)$
In particular from the above inequality we can conclude that $\lim_{a\rightarrow M^{-}}R(a)=0$ that is a contradiction. The conclusion of Lemma 4.3 is now the same as in Lemma 3.3. ∎
4.3. When $\left|\Sigma\right|>\left|\mathbb{S}^{2}\right|$: a sequence of weight functions
Due to the volume bound $|\Sigma|\geq|\mathbb{S}^{2}|$, we cannot directly compare $w_{1}(G)$ with $w_{1}(G^{\star})$ when $\left|\Sigma\right|>\left|\mathbb{S}^{2}\right|$ simply because $G,G^{\star}$ are defined in different intervals: $(0,M)$ and $(0,4\pi)$, respectively, with $4\pi<M$ (recall, $M=|\Sigma|$). Moreover $M=4\pi$ if and only if $\Sigma=|\mathbb{S}^{2}|$.
We define a sequence of weights $G^{\star}_{\varepsilon}$ as follows. Let $\varepsilon_{0}>0$ be a sufficiently small number to be fixed later, and let $\varepsilon\in(0,\varepsilon_{0})$. Let
@@
and moreover $G>0$ in $(0,M)$. Then there exists $\varepsilon_{0}>0$ such that for all $\varepsilon\in(0,\varepsilon_{0})$, we have $G\geq G^{\star}_{\varepsilon}$, where
$G^{\star}_{\varepsilon}=\begin{cases}G^{\star}\,,&{\rm in\ }(0,a_{\varepsilon})\\ \varepsilon&{\rm in\ }(a_{\varepsilon},M-b_{\varepsilon})\\ G^{\star}(a-M)\,,&{\rm in\ }(M-b_{\varepsilon},M).\end{cases}$
See Figure 1. The function $G^{\star}_{\varepsilon}$ is continuous on $(0,M)$ and converges uniformly to $G^{\star}$ in $(0,4\pi)$ and to 0 in $(4\pi,M)$.
Figure 1. the function $G^{\star}_{\varepsilon}$
@@
From Lemma 4.3 we deduce that, since $G\geq G^{\star}_{\varepsilon}$ for all $\varepsilon\in(0,\varepsilon_{0})$ we see:
(4.7) $w_{1}(G)\leq w_{1}(G^{\star}_{\varepsilon}).$
Now, the upper bound (2.2) in Theorem 2.4 is a consequence of the following lemma and the fact that $\mu_{1}(\mathbb{S}^{2},A^{\star})=\nu(\nu+1)$ if $\nu\in(0,\frac{1}{2}]$. The explicit expression of $\mu_{1}(\mathbb{S}^{2},A^{\star})$ is proved in Appendix A. Recall that by $\lambda_{1}(\Omega,A)$ we denote the first magnetic Dirichlet eigenvalue on $\Omega$ with potential $A$, see (1.1).
@@
    $w_{1}(G^{\star}_{\varepsilon})\leq\lambda_{1}(D_{a_{\varepsilon}},A^{\star})$, where $\lambda_{1}(D_{a_{\varepsilon}},A^{\star})$ is the first magnetic Dirichlet eigenvalue of a geodesic disk $D_{a_{\varepsilon}}\subset\mathbb{S}^{2}$ of area $a_{\varepsilon}$ with center $x_{0}$ and with $A^{\star}=-2\pi\nu\star d\psi_{x_{0}}=\nu d\theta$
@@
(4.8) $w_{1}(G^{\star}_{\varepsilon})=\min_{0\neq f\in\mathcal{F}}\frac{\int_{0}^{M}\left(G^{\star}_{\varepsilon}(a)f^{\prime}(a)^{2}+\frac{4\pi^{2}\nu^{2}}{G^{\star}_{\varepsilon}(a)}f(a)^{2}\right)da}{\int_{0}^{M}f^{2}(a)da}\\ \leq\min_{\begin{subarray}{c}0\neq f\in\mathcal{F}\\ f\equiv 0{\rm\ on\ }(a_{\varepsilon},M)\end{subarray}}\frac{\int_{0}^{M}\left(G^{\star}_{\varepsilon}(a)f^{\prime}(a)^{2}+\frac{4\pi^{2}\nu^{2}}{G^{\star}_{\varepsilon}(a)}f(a)^{2}\right)da}{\int_{0}^{M}f^{2}(a)da}\\ =\min_{\begin{subarray}{c}0\neq f\in\mathcal{F}\\ f(a_{\varepsilon})=0\end{subarray}}\frac{\int_{0}^{a_{\varepsilon}}\left(G^{\star}(a)f^{\prime}(a)^{2}+\frac{4\pi^{2}\nu^{2}}{G^{\star}(a)}f(a)^{2}\right)da}{\int_{0}^{a_{\varepsilon}}f^{2}(a)da}=\lambda_{1}(D_{a_{\varepsilon}},A^{\star}).$
The last equality holds since for the disk with Aharonov-Bohm potential with pole in the center and flux $\nu\in(0,\frac{1}{2}]$, there exists a first real and radial eigenfunction, as in the Neumann case. This proves i). Point ii) is proved in Lemma A.3. Namely, in Lemma A.3 we show that the first magnetic Dirichlet eigenvalue on a spherical cap with pole at its center coverges to the first eigenvalue on the sphere with two opposite poles when the cap invades the whole sphere. ∎
@@
(4.9) $\mu_{1}(\mathbb{S}^{2},A^{\star})=\mu_{1}(\Sigma,A)\leq w_{1}(G)\leq w_{1}(G^{\star}_{\varepsilon_{1}})<w_{1}(G^{\star}_{\varepsilon_{2}})\\ <\cdots<w_{1}(G^{\star}_{\varepsilon_{n}})\leq\lambda_{1}(D_{a_{\varepsilon_{n}}},A^{\star})\to\mu_{1}(\mathbb{S}^{2},A^{\star}).$
By Lemma 4.3 we have that the inequalities $w_{1}(G^{\star}_{\varepsilon_{n}})<w_{1}(G^{\star}_{\varepsilon_{n+1}})$ are strict, since by construction $G^{\star}_{\varepsilon_{n}}>G^{\star}_{\varepsilon_{n+1}}$ on a set of positive measure. Hence all the inequalities must be equalities, which is not possible, unless $M=4\pi$ and hence $\Sigma=\mathbb{S}^{2}$. Thus we are in the case of the sphere $\mathbb{S}^{2}$ with two poles $x_{0},y$ as in Corollary 2.5. If equality holds, then $G=G^{\star}$ in $(0,4\pi)$ and this implies that $y=-x_{0}$ by Lemma 3.2.
@@
The proof of Theorem 2.3 is exactly the same as that of Theorem 2.4, except that now $G^{\star}_{\varepsilon}$ is defined as
$G^{\star}_{\varepsilon}=\begin{cases}G^{\star}\,,&{\rm in\ }(0,a_{\varepsilon})\\ \varepsilon\,,&{\rm in\ }(a_{\varepsilon},M).\end{cases}$
for all positive $\varepsilon\leq\varepsilon_{0}$, for a suitable $\varepsilon_{0}>0$. The equality case is discussed as in Theorem 2.4.
4.7. Hersch’s Theorem fails: sharpness of Theorem 2.4
In this section we want to show that without an assumption on the Gaussian curvature, we cannot have a result like Hersch’s Theorem for the usual Laplace-Beltrami operator.
Theorem 4.5.
One has
$\sup_{\Sigma,\,z_{1},z_{2}\in\Sigma}|\Sigma|\,\mu_{1}(\Sigma,A^{(\nu)}_{z_{1},z_{2}})$$\infty$
for all $\nu\notin\mathbb{Z}$.
Proof.
We prove the theorem with the following family of surfaces. For any $L>0$ let $\Sigma_{L}$ be the cigar-like surface obtained by gluing two hemispheres $\Omega^{+}$, $\Omega^{-}$ of radius $1$ to the boundary components of the cylinder $\Gamma_{L}=\{(x,y,z)\in\mathbb{R}^{3}:x^{2}+y^{2}=1,0\leq z\leq L\}$, see Figure 2. The surface $\Sigma_{L}$ is punctured at the poles $p^{+},p^{-}$ of the hemispheres. Let $A^{(\nu)}_{p^{+},p^{-}}$ be the Aharonov-Bohm potential with poles $p^{+},p^{-}$ and flux $\nu$. By gauge invariance, we can take $\nu\in(0,\frac{1}{2}]$. We show that
(4.10) $\mu_{1}(\Sigma_{L},A^{(\nu)}_{p^{+},p^{-}})\geq\nu^{2},$
independently on $L$. In fact, from Appendix A we have that $\mu_{1}(\Omega^{\pm},A^{(\nu)}_{p^{+},p^{-}})=\nu(\nu+1)$, while from Appendix B we have that $\mu_{1}(\Gamma_{L},A^{(\nu)}_{p^{+},p^{-}})=\nu^{2}$. To simplify the notation, for the rest of the proof we shall write $A=A^{(\nu)}_{p^{+},p^{-}}$. Then for all $u\in H^{1}_{A}(\Sigma_{L})$ we have
$\int_{\Gamma_{L}}|d^{A}u|^{2}\geq\nu^{2}\int_{\Gamma}|u|^{2}$
and
$\int_{\Omega^{\pm}}|d^{A}u|^{2}\geq\nu(\nu+1)\int_{\Omega^{\pm}}|u|^{2}.$
Summing up, since $\nu(\nu+1)\geq\nu^{2}$, we get
$\int_{\Sigma_{L}}|d^{A}u|^{2}\geq\nu^{2}\int_{\Sigma_{L}}|u|^{2}$
for any $u\in H^{1}_{A}(\Sigma_{L})$, hence (4.10). It follows that
$\left|\Sigma_{L}\right|\mu_{1}(\Sigma_{L},A^{(\nu)}_{p^{+},p^{-}})\geq 2\pi\nu^{2}(L+2)$
which grows linearly to infinity as $L\to\infty$. The assertion follows. ∎
Figure 2. The cigar-like surface $\Sigma_{L}$
A final remark is that in Theorem 4.5 we give examples of closed surfaces for which we have lower bounds of the type $\nu^{2}K$ which are independent on the area, where $K$ is an upper bound on the Gaussian curvature ($K=\frac{1}{R^{2}}$ for $\Sigma_{L,R}$ in the proof of Theorem 4.5).
@@
It is enough to consider the example in the previous subsection and remove one of the two spherical caps $\Omega^{+}$ or $\Omega^{-}$.
Appendix A Spectrum of the punctured sphere, and more
A.1. Spectrum of the punctured sphere